OpenStudy (anonymous):

find the smallest possible integer value of n for which 99n is a multiple of 24.

OpenStudy (amistre64):

you want to find the least common multiple of 99 and 24 then

OpenStudy (amistre64):

do you know how to factor the numbers into smaller amounts?

OpenStudy (anonymous):

LCM

OpenStudy (amistre64):

a property you can use is also \[\frac{ab}{gcd(a,b)} = lcm(a,b)\] but that tends to be more advanced

OpenStudy (anonymous):

yeah, I know.

OpenStudy (anonymous):

I know the answer, but I don't the working.

OpenStudy (anonymous):

I dont know the working.

OpenStudy (amistre64):

hmm, it works upon the factorizations of 99 and 24

OpenStudy (amistre64):

do you know about prime numbers?

OpenStudy (anonymous):

yeah, I do.

OpenStudy (amistre64):

all integers have a unique prime factorization of integers. This helps us to know what they have in common, and what they would need as well

OpenStudy (amistre64):

what is the prime factorization of primes for 99?

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

3x3x11

OpenStudy (amistre64):

good; and the prime factorization of 24?

OpenStudy (anonymous):

2x2x2x3

OpenStudy (amistre64):

excellent; now the next step has many methods that can be used; i perfer this one: stack the factorizations one aboove the other such that like factors line up 99: 3 3 11 24: 2 2 2 3 ---------------- LCM: then drop down all the columns to fill in the LCM 99: 3 3 11 24: 2 2 2 3 ---------------- LCM: 2 2 2 3 3 11 , now multiply all those values together, what do we get?

OpenStudy (anonymous):

792?

OpenStudy (amistre64):

792 is correct; this gives us the value that is the smallest multiple that 99 and 24 have in common therefore: 99*n = 792 n = 792/99 , what value do we get for "n" now?

OpenStudy (anonymous):

n = 8

OpenStudy (amistre64):

correct. now, that is only the method I use; there are many other ways to approach this

OpenStudy (anonymous):

oh, thanks. really appreciate it.

OpenStudy (amistre64):

youre welcome, and good luck ;)